OpenStudy (anonymous):

x+1=9x^3+9x^2 im supposed to solve the polynomial by factoring and then using the zero product principle

OpenStudy (dumbcow):

Factor by grouping (9x^3 +9x^2) - (x+1) = 0 9x^2 (x+1) -(x+1) = 0 (x+1)(9x^2 -1) = 0 --> x+1 = 0 x = -1 --> 9x^2 -1 = 0 x = +-1/3

OpenStudy (anonymous):

thanks

OpenStudy (dumbcow):

yw